What is the Molarity of a solution that contains 5.85 grams of NaCl in 0.50L of a water solution?




1
Expert's answer
2021-10-27T06:41:07-0400

Molar mass of NaCl= 58.44

5.85/58.44= 0.1mol

Molarity= 0.1/0.5= 0.2mol/L